Calculate the mass of ferric oxide (FeO3) that will be produced by complete oxidation of 100g of Iron
1
Expert's answer
2020-07-09T14:54:34-0400

Solution:

2Fe+3O2=2FeO3

n(Fe)=m(Fe)/M(Fe)=100/56=1.786 mol.

n(FeO3)=n(Fe)=1.786mol.

m(FeO3)=n(FeO3)*M(FeO3)=1.786mol*104g/mol=185.7g.

Answer: 185.7g